john can type a 40, page document in 24 minutes, susan can type it in 30 minutes, jack can type it in 24. working together, how much time will it take them to type the same document?
Answer by checkley77(12844)

1/40+1/24+1/30=1/X
(3+5+4)/120=1/X
12/120=1/X
1/10=1/X CROSS MULTIPLY.
X=10 MINUTES ANS.